OpenStudy (anonymous):

The solution to 2x - 2 + 5 = 13 is _____. 10 8 5 3

OpenStudy (uri):

2x=13-5+2 2x=10 x=5

OpenStudy (jh3power):

it's same as 2x+3=13 2x+3-3=13-3 2x=10 x=5

OpenStudy (jh3power):

2x/2=10/2..